THE ELECTRONIC STRUCTURE OF THE PLUTONYL ION

\underline{JEAN-PHILIPPE BLAUDEAU}, BRUCE E. BURSTEN and RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Restricted Hartree-Fock (RHF) and spin-orbit configuration-interaction (SOCI) calculations were performed on the ground and low-lying excited states of the plutonyl ion, PuO_{2}^{2+}. These results are compared to density functional theory (DFT) results obtained using generalized gradient corrections. The low energy transitions are f→f. Information on the vibrational modes will also be presented. For the ab initio work, the actinides are modeled with relativistic effective core potentials and Gaussian correlation consistent double-zeta plus polarization basis sets. The DFT work used Slater basis sets.